### v1

# Scaling the Authoring of AutoTutors with Large Language Models

Abstract.
Large Language Models (LLMs) have found several use cases in education, ranging from automatic question generation to essay evaluation. In this paper, we explore the potential of using Large Language Models (LLMs) to author Intelligent Tutoring Systems. A common pitfall of LLMs is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees. We posit that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results. Based on this principle, we create a sample end-to-end tutoring system named MWPTutor, which uses LLMs to fill in the state space of a pre-defined finite state transducer. This approach retains the structure and the pedagogy of traditional tutoring systems that has been developed over the years by learning scientists but brings in additional flexibility of LLM-based approaches. Through a human evaluation study on two datasets based on math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4. MWPTutor is completely modular and opens up the scope for the community to improve its performance by improving individual modules or using different teaching strategies that it can follow.
MWPTutor
I see your solution. You seem to have made some errors. How about we start fresh and do it step by step?
Okay, I’m not sure where to start though.
What should be the first step?
I think Cody’s starting amount of tickets, which is 49, should be the first step.
Tell me this instead: How many tickets does Cody have left after spending 25 on a beanie from his original 49 tickets?
Oh, he would have 24 tickets left.
24 is correct, well done. What is the next step?
Next, he won 6 more tickets, so I think we add that to the 24 he already has.
Okay, can you do that calculation?
Yes, 24 + 6 is 30, so Cody would have 30 tickets in total.
30 is the correct answer to the problem. Well done!!! Don’t hesitate to contack me again if you have any troubles.
GPT4 Tutor
Hi! Could you walk me through your solution?
I’m not sure where to start. I know Cody won 49 tickets and spent 25 on a beanie, but I don’t know how to figure out how many he has left.
That’s a great starting point. So, if Cody had 49 tickets and he spends some on a beanie, what kind of math operation would you use to find out how many tickets he would have left?
Subtraction, I think.
Exactly, subtraction is the right operation. Can you try subtracting the number of tickets Cody spent on the beanie from the number of tickets he initially won? What do you get?
— 12 more tutor utterances —
28 tickets.
That’s correct. Now, if we subtract another ticket from the 28, what number do we get?
27 tickets.
Yes, that’s right. And another ticket taken from 27? What does that leave us with?
26 tickets.
Exactly. And if we take one more away from 26, where does that get us?
25 tickets.
— 13 more tutor utterances —
1. Introduction
Computer-based automated tutoring has long been seen as the method to scale up the benefits of expert human tutoring to millions of students, a problem originally posed by Bloom (Bloom, 1984). Of various attempts to do so, the AutoTutor family (Nye et al., 2014) is perhaps one of the more notable ones. This family of tutors simulates natural language conversations with students based on pre-defined scripts, aiming to emulate one-on-one tutoring. While the AutoTutor family has shown substantial learning gains (Nye et al., 2014), it is difficult to expand it to different topics. Each expansion requires a lot of expert labor to incorporate the required domain knowledge for every concept in the curriculum and appropriate teaching strategies (Koedinger et al., 2004).
Recent developments in Large Language Models (LLMs), such as GPT (OpenAI, 2023), or FlanT5 (Chung et al., 2022), have opened up a new venue for computer-based tutoring where the LLMs play the role of the teacher. While previous work has shown the success of LLMs for tutoring (Rajala et al., 2023; Baidoo-Anu and Ansah, 2023) the non-modular black-box nature of these models makes it impossible to make incremental improvements or control the pedagogy it pursues in assisting the student. The uninterpretable and uncontrollable nature of LLMs makes it difficult for experts or teachers to make changes in their strategy. It is also difficult to develop trust towards such models, not least because they might be open to jailbreak attacks (Liu et al., 2023) as well as privacy concerns. Finally, the proprietary nature of highly performant LLMs means that the tutor needs a persistent internet connection and an OpenAI account to function, and we are highly dependent on OpenAI policies. This could be bypassed by fine-tuning a local open LLM, but doing so leads to either prohibitively high resource requirements, or significantly worse performance (see FlanT5 vs ChatGPT in (Macina et al., 2023)).
The design process of AutoTutor and other similar intelligent tutoring systems can be split into two distinct stages, carried out by different, yet potentially overlapping, sets of stakeholders. The first stage is the general framework level design, carried out mostly by learning scientists (possibly along with computer scientists). This includes designing the pedagogy, the mechanisms required to interact with the student, and potentially even authoring tools. This is typically done only once for an entire domain. The second stage, performed by either teachers or domain experts, is to author the scripts111see Appendix A in (Graesser et al., 2000) for an example script, which needs to be done separately for every topic, question, or concept. In the standard framework, both these stages are done by humans. In contrast, in the LLM-only framework discussed in the previous paragraph, both these stages are carried out (implicitly or explicitly) by LLMs.
In this paper, we consider a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts. This idea has previously been hinted at by Cai et al. (Cai et al., 2019), who propose using a question-answering system to automatically list potential misconceptions for a given topic. We extend this idea to have an LLM author the entire script for the tutoring system, which is accurate enough to be deployed without human validation but is fully open to human validation.
Our contributions in this paper are as follows:
We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for a Large Language Model to author it.
We design MWPTutor, a tutor for math word problems based on the above principles.
We compare the performance of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
2. Background
Tutoring is an integral part of teaching where the teacher attempts to identify confusions or misconceptions a student might have, and rectify them. This is done in a way to minimize their future re-occurrence. Bloom et al. (Bloom, 1984) showed that tutoring is extremely important for education, as a student receiving one-on-one tutoring can perform much better than a student only getting lectures. As opposed to lecturing, tutoring is more personalized, and therefore only feasible in smaller groups. This lower teacher-to-student ratio needs to be balanced by a higher number of tutors, which is expensive and not always feasible.
2.1. Rule-Based Systems
To combat the above scarcity, Intelligent Tutoring Systems (ITSs) were already being developed since the early 1980s (Sleeman and Brown, 1982). The idea was to roleplay a human tutor with a computer program that could interact with the student and fulfill the requirement of personalized tutoring. The computer program may not fully replace human tutors but often provides an affordable and scalable addition. An important aspect that differentiates ITSs from simply showing the students the problem solution is their ability to provide step-by-step feedback which promotes long-term learning as opposed to short-term problem-solving success. This was often implemented using rules that check for certain errors using string-matching and act on them with pre-defined templates. For example, LISP Tutor (Anderson et al., 1985) contains rules to account for common programming errors made by students.
With advancements in AI, ITSs transitioned to dialog as a method of tutoring, moving ITSs from being information-delivery systems to dialog-based agents. One of the key players in this field was the first version of AutoTutor (Graesser et al., 2001), which replaced the rules in traditional ITSs with a full finite state transducer which they called the Dialog Advancement Network. This allows the students to interact in natural language and the tutor processes their inputs using methods like latent semantic analysis and attempts to respond using a set of predefined rules. Over the next two decades, various tutors such as the AutoTutor explored different domains, tutoring strategies, affective statements, etc. to become one of the most comprehensive ITSs (see (Nye et al., 2014) for the summary of developments).
With concepts and tutoring strategies getting more complex, writing the rules governing the finite state transducers quickly becomes time-consuming. Not knowing the student misconceptions in advance, domain experts have to come up with rules for numerous possibilities, introducing a different set of scaling issues. One hour of interaction with a student costs over 100 hours of domain expert time to author the ITS (Koedinger et al., 2004). This is in addition to the time spent by learning scientists designing the pedagogy and dialog strategies. However, this is usually done only once for the topics in a subject area. As a result, this limits the AutoTutor to a small set of subject areas, restricting their impact. Nevertheless, there are many works in the space of rule-based ITS (Paladines and Ramirez, 2020), which have been reasonably successful in their domains.
2.2. Large Language Models
Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2022). With recent improvements in LLMs and the availability of LLM-as-a-service models like GPT (OpenAI, 2023) and BARD (Yarkin and Coon, 2022), it has become easy to use LLMs as tutors by prompting them with the desired qualities of a tutor and then letting them interact with a student. Some works in the programming domain like GPTutor (Chen et al., 2023) and TestGPT (Phung et al., 2023) attempt to use GPT4 similar to information delivery systems. There have also been systems designed to function as Dialog agents, like Ruffle&Riley (Schmucker et al., 2023) which includes multiple GPT4 agents functioning with different prompts, and SPOCK (Sonkar et al., 2023) which attempts to distill out tutoring knowledge from GPT4 into a smaller LLM. A notable commercial usage of a LLM Tutor is seen in Khanmingo,222khanacademy.org/khan-labs which uses the GPT4 backend for an interactive tutor.
While LLM tutors solve the scaling issues, it introduces a whole new set of complications, primarily related to its probabilistic output and proprietary models. The behavior of GPT4 can only be controlled by making changes to its initial prompt or by inserting prompts in between, making it hard to introduce the nuanced pedagogical strategies developed by learning scientists over the last two decades. Even for simpler strategies, its probabilistic nature leaves it open to being exploited by users who behave in unwanted ways, often referred to as jailbreak attacks (Liu et al., 2023). Even for benign users, GPT4 can often hallucinate incorrect information and output it with confidence (Zhang et al., 2023; Cui et al., 2023) which may adversely affect the student. All this, in addition to their hard-to-understand nature, leads to a distrust of GPT4 among educators and policymakers, causing it to be banned in several US public schools.333bestcolleges.com/news/schools-colleges-banned-chat-gpt-similar-ai-tools
2.3. Our work
The main benefit of the fully human-authored setting comes from the fact that it is easy to control, easy to interpret, and therefore easy to trust. Most of these benefits come from the pedagogy design stage, and not from the work of the domain experts. The LLM-based approaches, on the other hand, mostly speed up the domain experts’ tasks. To bring together the best of both worlds, we propose an expert-designed pedagogy to be assisted by an LLM in the dialogue part. The rest of this paper is structured as follows: Section 3 introduces our general design which applies to several domains, Section 4 describes the specific design of MWPTutor, Section 5 quantitatively evaluates MWPTutor, and Section 6 looks qualitatively at weaknesses of both GPT4 and MWPTutor.
3. Design Principles
While this work mainly focuses on Math Word Problems tutoring, our broader goal is to present a paradigm that can be used across different subjects. Before the math-specific aspects, we first describe the model design principles. The output is the implementation of a pedagogical goal as a template state space for a finite state transducer similar to dialog advancement networks, which can then be filled by an LLM. For the pedagogical strategy, a good doctrine to follow can be Get the student to the correct answer. If possible, do so without telling them the answer outright
A finite state transducer is characterized by the combination of state space and a transition function. The state space can be seen as the memory of the finite state transducer, which stores one or more variables, each of which can take two or more values. In our case, we have two variables, the solution step and the strategy, which factorizes our state space into two subspaces, detailed in Section 3.1 and Section 3.2 respectively. The transition function specifies the next stage given the current state and the (student) input and is highly domain-dependent, but we provide examples as they come up alongside the state space descriptions
First, consider a math word problem and a history question to illustrate the possible range of topics which a tutor might apply.
The optimal tutoring support differs between the two examples. While the math word problem starts with subquestions to compute intermediate quantities, answering the history questions first requires providing historical context.
3.1. Solution Decomposition
To begin creating the state space for a finite state transducer, we first need to list the solutions for the problem, creating the solution step space. As a first design principle, we want to have the solution broken down into multiple parts. For the math word problem example, this could look like a tree:
To complete the question, the student needs to perform A1 followed by A2. Alternatively, they can complete B1 and B2 in any order and then B3. We transition out of a step if the student arrives at the right-hand side of the equation in the (sub)question. These dependency structures could look very different based on the domain, like for the history question, for example:
In this case, a correct answer requires the student to complete any 4 out of C1-C6 in any particular order to be considered successful. The transition function here is more complicated, but one way would be to check if all underlined key terms have been mentioned.
Splitting a problem into smaller parts is beneficial for both human learning (Rich et al., 2018; Song and Becker, 2014) and for LLM solvers (Dhuliawala et al., 2023; Nye et al., 2021). Getting an LLM to split the problem into smaller parts allows us to thereafter focus the LLM on smaller parts of the solution, which can also reduce context lengths. In AutoTutor terminology, these solution steps would correspond to “expectations”.
3.2. Pedagogical Strategy
Having broken down a problem into steps, we now need to have a strategy to help the student achieve each step. The general idea is to start vague, hoping for the student to do as much of the work as possible, but then keep getting more specific until the student can complete the step. For the math word problem example, a toy strategy could be:
This set of pedagogical strategies forms our strategy space. We can transition from D1 to D2 and then to D3 if the conditions for transitioning to the next solution step are not met, at which point we loop in D3. For the history example, we have:
The actual set of moves can be quite complex. For P2, we would need separate utterances for every possible combination of missing keywords. Ideally, these moves should be designed by learning scientists as was the case for AutoTutor.
The overall state space is the cross-product of the solution state space and the strategy space. Figure 2 shows what the full state space of our MWP example could look like. For each element of the strategy space, we would require prompt templates that can be filled with the solution step and passed on to LLMs to generate the actual utterance. Running each prompt for every step in the solution space maps every element in the state space to an utterance. Not all utterances need to be unique, and some, like P1 above, might not necessitate an LLM to be generated.
3.3. Scaffolding the LLM
While modern LLMs fine-tuned with RLHF are quite good at following instructions, they can at times struggle with more complex tasks. Because of the probabilistic nature of LLMs, reattempting the same problem can lead to different answers. To take the best advantage of these facts, we want to have automated checks for every LLM turn in the strategy space, so calls can be repeated until some checks have been met. For example, D3 could require that the numerical answer (450 for A1) is included in the utterance, while P2 could require the missing keywords to be absent. These checks can take different forms depending on the domain, like limiting the length of an utterance or avoiding the use of more complex language.
3.4. Control and Reliability
One of the primary goals of having LLMs author a state space rather than having them interact directly with the students is that in this way, one can examine and trust the model without having to necessarily trust the LLM used to generate it. Based on this, we store and make it possible to inspect beforehand the entire utterance space. Changes to the utterances themselves can be made on-the-go, e.g. if a teacher feels it is better to split our example in P2 into two sentences, they can simply edit it to say “Who was assassinated?” and “Who carried out the assassination” with no further changes required. Changes to the solution space, however, need the utterances to be regenerated.
Caching the utterance LLMs like GPT also makes the system immune to changes made to GPT in terms of availability, cost, or rate limits. For open LLMs, this allows for the use of Larger LLMs which not every user can run locally. The most important benefit, however, is the potential for teachers/domain experts to edit the state space should they want, therefore allowing us to combine human expertise with machine efficiency. Live calls to an LLM can often help and can be incorporated into the state space. For example, in the MWP we could make a call to GPT to take over a step in between D2 and D3 if we want to minimize the chance of D3. Should live calls be implemented, however, there should ideally be non-LLM fallbacks ie we should have cached and/or procedural alternatives available should a call to an LLM not be possible at deployment time.
4. MWPTutor
We now detail the design process of our system MWPTutor, an instantiation of the above principles. MWPTutor is adapted to help students solve math word problems. The scenario, originally defined by MathDial (Macina et al., 2023), involves a student approaching a tutor after unsuccessfully attempting to solve a math word problem. The student’s incorrect solution is provided as a part of the input. GSM8k (Cobbe et al., 2021), the source for MathDial’s problems includes a single step-by-step solution to each problem, which can be seen as a fulfillment of the solution step space. However, to not become constrained to the availability of such solutions, and to allow diverse solution paths, we restrict MWPTutor to only the final numerical answer to the problem, and generate the solutions ourselves.
symbol indicates student utterance inputs, while the
symbol indicates model output. The solution step space is collapsed for clarity
4.1. Solution Decomposition
We instantiate the solution steps as short pieces of text describing a step in the math word problem, including a single mathematical calculation. We generate multiple solution paths with our backend LLM, with each step needing to contain a single calculation demarcated in double angular braces (e.g. , see Table 3). This requirement standardizes the information density of a step and also keeps things consistent with GSM8K which uses a similar notation. It also allows us to check for calculation errors and discard erroneous paths. Once multiple solution paths have been obtained, we discard the paths not leading to the final correct solution, and group the rest together using exact string match on the calculations into a “tree” of solution steps. Our running example, except for the angular brackets, fits this criterion as a tree with two branches at the root and no branches elsewhere.
4.2. Solution Alignment
Although not a part of design principles, we wish to not make the students repeat a step they have already performed correctly. To achieve this, we match steps in the original student solution with the closest path in the solution step space. While AutoTutor achieves this using latent semantic analysis and string-matching techniques, we found that modern counterparts, such as BERTScore (Zhang et al., 2020) are surprisingly not an apt replacement, perhaps because GPT4 outputs very long step descriptions, and BERT does not weigh and compare numbers and mathematical operations well.
MWPTutor currently assumes that the student will present their solution with one step in each line. For each student step, we check if the right-hand side of any of the possible follow-up steps is present in it. If more than one follow-up matches, we pick the one with the highest BERTScore. We then proceed down that node of the solution tree and the next student step. If no follow-up matches, the current step is considered to be the first point of error. The pseudo-code for this is shown in Algorithm 1.
Input: gold and student solutions.
Output: aligned solution step, if exists
Filter out next solution steps whose RHS is in the current student step
Move to next step
4.3. Pedagogical Strategy
To instantiate our pedagogy, we take inspiration from AutoTutor, where the primary sequence of steps is a pump→hint→prompt→ assertion (Section 2 of (Graesser, 2016)). Graesser et al. in their work Teaching Tactics in Autotutor (Graesser et al., 2000) define hints as “questions that lead the student to the desired information”, and assertions (which they call elaborations in this work) as “succinctly assert the desired information.” For prompts, they use fill-in-the-blanks type utterances leaving out a keyword. In our case, the “keyword” is the equation, so instead we explain the calculation to be performed in a natural language. These are preceded by a “pump”, which is the generic utterance “What should be the next step?” The overall strategy then proceeds as follows:
| Pump |
asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
| Hint |
is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose. It essentially gives out the step and asks the student for the calculation. This utterance should take the form of a question, and the only way to correctly answer the question should be to successfully execute the step. |
| Prompt |
is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
| Assertion |
is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
In our running example, hints, prompts, and assertions would correspond to D1, D2, and D3 respectively. The pump is not included.
For scaffolding the LLM, both hints and prompts are required to avoid explicitly outputting the right-hand side of the equation in the step, with the hint additionally having to avoid saying any words that indicate arithmetic operators, currently “add, subtract, sum, multiply, product, divide, quotient.” The assertion, on the other hand, must contain the right-hand side. For pumps, there is no use of an LLM, so we do not need a check.
4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to be able to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
The fact that we currently ignore all non-numeric content in the student response is one of the major limitations in our work, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
Inputs: CurrentSteps from Solution step space, and Student Utterance
Output: Degree of the match, Best matching step
CurrentSteps is a singleton except for “pump.”
Go to next solution step
4.5. Putting it All Together
Figure 3 summarises the overall state space of MWPTutor. It can be seen as a simplified version of the state space of AutoTutor (see Figure 1 in (Cai et al., 2019)). The expectation matching is implicitly carried out by the solution aligner and the pump state. To obtain to full finite state transducer structure, one must roll out the states along the solution space. Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs, where is the number of steps.
5. Experiments
In this section, we the performance of MWPTutor and GPT4 in terms of how well they can follow our previously mentioned doctrine. We test both models in the interactive setting proposed by Macina et al. (Macina et al., 2023) which allows us to safely compare imperfect models without having a negative impact on human users.
5.1. Setup
5.1.1. Setting.
The interactive tutoring setting was proposed in MathDial as a faster, cheaper, and safer alternative to human experiments, where instead of interacting with human students, a Tutoring System interacts with another LLM role-playing a student. In their data collection experiments, MathDial found that most teachers believed that the student LLM behaved very much like an actual student in most cases. We adopt their setting as is for our experiment section.
Note that for an LLM, the interactive tutoring setting is somewhat related to the MWP solver setting often used as a metric to measure the reasoning abilities of said LLM, but there are significant differences. Unlike the MWP solver setting, the final answer is already provided, so the LLM does not need to come up with it. The difficulty arises from the fact that this LLM needs to get the student LLM to get to this answer, and the student LLM is prompted to not budge from its starting wrong answer without good reason. Despite these differences, some skills like Math Reasoning Abilities transfer between the two settings, so we can expect the final performances in both settings to correlate with each other.
5.1.2. Models.
At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for MWPTutor and as the baseline single prompt LLM. To best align the single prompt GPT4 with the MWPTutor doctrine, we ask it to roleplay a tutor, setting the primary objective of getting to the correct answer as the goal of the conversation, and the secondary objective of not revealing the answer as a personality trait. Further, to test the potential of an open-source LLM as a backend, we have a second version of MWPTutor with a MetaMath-13B (Yu et al., 2023). MetaMath is a version of Meta AI’s open source LLM LLaMa666https://llama.meta.com/ fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems. We refer to these three models as MWPTutor, GPT4, and MWPTutor respectively.
5.1.3. Datasets.
To implement MWPTutor, we require a dataset with math word problems and their solutions, as well as incorrect student solutions. We make use of the test sets of two datasets, namely MathDial (Macina et al., 2023) and MultiArith. MultiArith consists of simple Math Word Problems that are usually solvable in 1 to 3 steps. GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2023), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting. With this dataset, we want to test if we can achieve the same performance in the interactive tutoring setting. This dataset does not provide any wrong solutions needed to instantiate the student model, so we make use of errors made by a MetaMath solver as the original student solution where possible, and leave the remaining questions as “No idea how to begin”.
MathDial consists of problems from the pre-existing GSM8k (Cobbe et al., 2021). Problems in GSM8k are meant to be solvable by a bright middle school student using only basic arithmetic operators (, , , ), taking up to steps to solve. Of these, MathDial picks problems that have official solutions taking steps or fewer, that GPT3.5 consistently failed to solve. This makes MathDial a significantly harder dataset to teach, and results on this dataset are expected to show the differences between the models more clearly.
5.1.4. Metrics.
Keeping in line with our doctrine, our primary metric of evaluation is the Success Rate (S), which is measured as the fraction of conversations where the student says the correct answer eventually. The secondary metric is Telling Rate (T), which is the fraction of conversations where the tutor did not reveal the answer. The tutor is said to have revealed the answer when they say the final answer before the student has said it, and the student succeeds in getting to the answer thereafter. Both these metrics were originally defined by Macina et al. (Macina et al., 2023), but they do not offer a way of combining them into a single score. It is quite easy to increase success at the cost of increased telling. Therefore we also calculate the Adjusted Success (S-T) rate as the fraction where the student succeeded without having been told the answer and use the harmonic mean between this number and the success rate as an overall Tutoring Score (). We also offer utterance level evaluations in Section 5.3, but since we are unaware of the utterance level strategies from GPT4, we do it only for our model, and the metrics for the same are defined later.
5.2. Conversation Evaluation
We evaluate both success and telling for the MultiArith dataset by directly looking for the answer in student and tutor utterances respectively, and calculate the overall scores based on that. The top half of Table 1 summarises our results. All models are able to transfer their near-perfect solving scores to near-perfect tutoring scores. GPT4 performs slightly worse, although it has a telling rate. Both versions of MWPTutor have a perfect success rate. It must be noted that the metric is noisy since it only requires the number to match, with no restriction on the context, e.g. in the problem “How many groups of can be made out of boys and girls?” if the tutor mentions there are boys, we would count it as telling as the final answer happens to be . This means that the real rates could be somewhat lower for all models, and the difference between them is not significant. Nevertheless, such errors are unlikely to be frequent, and all models work reasonably well on these problems.
| MWPTutor | GPT4 | |||
| Metric | LLaMa | GPT | ||
MultiArith |
Success Rate | 1.00 | 1.00 | 0.98 |
| Telling Rate | 0.01 | 0.01 | 0.00 | |
| Adjusted Success | 0.99 | 0.99 | 0.98 | |
| Tutoring Score | 0.99 | 0.99 | 0.98 | |
MathDial |
Success Rate | 0.91 | 0.96 | 0.82 |
| Telling Rate | 0.33 | 0.11 | 0.00 | |
| Adjusted Success | 0.58 | 0.84 | 0.82 | |
| Tutoring Score | 0.71 | 0.90 | 0.82 | |
For the harder MathDial dataset, we wanted to avoid matching errors as described earlier, and so have human annotators mark out utterances where the student or the tutor says the final answer. Each conversation was annotated by annotators, and we took the majority in cases where they disagreed.
Table 1 summarises our result. While GPT4 maintains its 0% telling rate, MWPTutor gets more problems correct overall, and more when adjusted for telling. This gives MWPTutor an advantage on GPT4 in terms of the final score. MWPTutor also gets a higher success rate than GPT4, which is somewhat surprising since its backend model is significantly weaker than GPT4 in the MWP Solver setting (Yu et al., 2023). This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall, so we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason. One thing worth mentioning here is that once problematic questions have been identified for either version of MWPTutor, one can have human annotators redo only those questions, thereby increasing the success rate, something that is not possible for GPT4.
5.3. Utterance Evaluation
A major shortcoming of AI-on-AI-tutoring is that the AI student may not be affected by the sometimes subpar tutor responses. This is more common in MWPTutor than in GPT4, as the former can only pick from a pre-defined set of responses and not create new utterances on the fly in response to finer details in the student requirements. To understand if this is a major issue, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor for MathDial. Because the pumps are generic and very similar to each other and assertions essentially boil down to reading out the answers, we focus only on the hints and prompts in our evaluation.
Across the conversations, MWPTutor and the student model went over 135 solution steps. Of these went to the hint state, while reached the prompt state. We presented each of these utterances to human evaluators, with the past conversation, the original math word problem, and the current solution step as context. We boil down the requirements set on hints and prompts set in Section 4.3 to the following binary criteria which we asked the evaluators to evaluate the utterances on:
Grammatical: The utterance is linguistically and grammatically correct.
Not Reveal: The utterance does not reveal the final answer for the step.
Relevant (hints only): The hint clearly addresses the step, ie, the only way to correctly answer the question is to complete the current step.
Not Redundant (hints only): The hint is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
Specific (prompts only): The prompt is specific to the step, and not just a vague lead-on statement.
Helpful (prompts only): The prompt can potentially help the student make progress in the current step, given the previous student’s utterance.
The annotators were paid the Prolific suggested wage of 10 GBP/hour. We show the annotation user interface in Figure 6 and some example annotations in Table 2 and Table 3. Figures 4 and 5 summarise the responses from the evaluators. Almost all utterances are considered Grammatical by the evaluators, which is not surprising given they were generated by GPT4. Most of the utterances are also marked to be non-revealing, which is also not surprising given that any proposed utterances that directly revealed the answer were assigned to be regenerated. The evaluators further seem to agree that most hints were relevant to the step and that most prompts are specific, implying that they do what is appropriate given the current solution step.
The biggest potential weakness of MWPTutor comes in the form of its utterances making sense in the context of previous student utterances due to issues discussed in Section 4.4, which is where the final two metrics come in. In this regard, all prompts are considered to be helpful by a majority of evaluators which is a great sign. Hints, however, are found to be redundant in almost a third of cases, showing that handling of student’s responses to pumps could be better. The current heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion. Note that of the steps did not proceed to the hint state, implying more than half of the correct proposals by the student are being identified as such. Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a majority of correct rejections.
Overall these evaluations show that utterances generated for MWPTutor by the GPT4 backend are consistent with the strategy definitions, and the transition function uses them reasonably well.
6. Qualitative Assessment of Weaknesses
While both GPT4 and MWPTutor do quite well on quantitative metrics as seen above, these metrics do not fully grasp what it takes to make a good tutor. To better understand the preparedness of our models, we went over all conversations that happened on the MathDial dataset, for both GPT4 and MWPTutor and summarize our findings here.
Firstly, MWPTutor tends to sound monotonous and dry. Part of this is because of the pre-set format for utterances like all hints are hardcoded to begin with “Tell me this instead:,” but this can easily be resolved by having it sample from several rephrasings of the same utterance. More difficult to fix are the repetitive strategy and the lack of finer feedback or any affective statements, both of which are there by design. Introducing finer feedback requires a more fine-grained analysis of student utterances, which we found off-the-shelf NLP methods to be incapable of doing while maintaining a high degree of accuracy. Affective statements, like encouragements, challenges, etc. would further require accurate judgment of student confidence and correctness. In both these cases, we made a design choice to avoid the additional module, as they would create major failure nodes. GPT4 on the other hand, sounds a lot more natural, including finer feedback and probing, and even affective statements, although they are exclusively positive.
Further, since it matches correct answers with limited attention to context, MWPTutor can sometimes get false positives, and inadvertently reveal the answer. Consider the following excerpt:
Here, the student gives the wrong answer of , but MWPTutor mistakes the intermediate calculation of as the correct answer, and effectively reveals it to the student. GPT4 is able to understand context usually, although it is far from perfect. As we discuss later, GPT4 fails much worse on this question too.
While GPT4 conversations flow a lot better and sound human-like, it has a different set of issues. For one, GPT4 tends to overcomplicate questions or make redundant calculations. In one question, where the student failed to convert seconds into minutes, GPT4 gets the student to convert the seconds to hours first, and then back to minutes. This introduced precision errors, which it had to fix by rounding the final answer. In another question GPT4 has a student recalculate a value she had already calculated in a previous step. In yet another question GPT4 makes the student calculate the weekly rate of weight loss, when a simple multiplication of time period by 4 would have yielded the answer. The conversation in Figure 1 is also an example of where this happens in the simpler MultiArith dataset. Since it explores multiple paths and is also limited in the steps it can take, MWPTutor is less likely to run into this issue, and handles the first two cases mentioned here pretty well, though it makes the same error in the last case.
The biggest problem with GPT4, however, continues to be that GPT4 can lead the student to a wrong answer, sometimes even leading them away from the right answer. In all cases where GPT4 fails to get the student to the correct answer, it outright accepts a wrong answer as correct. Of MWPTutor’s failures one was a result of inadvertently revealing the final answer, while the other seems to have been marked as incorrect by the annotators because the tutor resorts to revealing rather quickly. In a further conversation, including the one regarding chicken feed described earlier, GPT4 misdirects the student onto the wrong path but somehow manages to get them to say the correct answer. Finally, in one question, GPT4 rejects a correct equation, and then goes off in a tangential direction, until the maximum limit of utterances is reached (this being the only instance where the utterance limit comes into play). It must be reiterated here that the final correct answer is given to GPT4 as part of its prompt. MWPTutor by design will never okay a wrong answer. The only place where it does go wrong, it accepts an answer of instead of , but this can be chalked down to the fact that the question was asking for the “difference” between two quantities, and did not make it clear whether the sign was required (the first quantity was smaller. Also GPT4 makes the same error. Both are marked as correct by annotators). It must, however, be noted that since it also relies on GPT4 for solution generation, MWPTutor can hallucinate, even if it did not happen in any of these conversations to the best of the authors’ judgement (We attempted to have the evaluators annotate the existence of wrong methods, but found too much disagreement in their responses to consider them viable). If this happens, however, one can fix them by changing the tree the first time they get noticed, something that is not possible with GPT4.
Overall, we conclude that despite good showing on metrics, both methods are still lacking in quality, and decide to not test them on real students until we can significantly improve them.
7. Discussion and Future work
We have stated throughout the paper that although self-sufficient, MWPTutor is still a prototype, and we explained in Section 6 that we still have a long way to go before making MWPTutor viable to be used on real students. In this section, we outline some directions for future work that could bring us closer to these goals.
7.1. Improving the Solution Step Space
The current solution step space takes the shape of a tree where solutions can only diverge from each step. But this is non-realistic as we know that all paths must lead to the same final answer and hence need to converge at some point. Further, the order of carrying out certain steps could be a partial order rather than a total order. This would currently be represented by listing out all possible orders as separate paths, but this causes a blowup in the number of steps. Adding a richer structure to the Solution graph could vastly reduce the size of the solution space.
In addition to this, each step of the solution should be made to follow a more standardized format. Having a more precise definition of the step would allow us to use LSA-like techniques to improve our correct answer detection. Pre-defining possible errors for each step would allow us to prepare to handle said errors. We can also have descriptions of each quantity in the LHS to give more pointed hints.
7.2. Improving the Strategy Space
The current strategy space is linear and does not allow for very targeted reactions to the student utterances or the original student solution. Once we have a better solution step space, we can look for different types of errors and diverge into different Prompts depending on the exact error should the student fail to get the answer correct after the Hint state. There can even be more than one Prompt revealing different amounts of information based on the complexity of the current step. Another thing entirely missing is the “Problematization” strategies described by Reiser (Reiser, 2004) and also used by MathDial. Alongside this, the matching requirements for solution alignment can be made stricter to better know where to start applying our intervention. This could also come with relaxing the format requirement of “one step per line” on the student solution. Both these would shorten the average lengths of the conversation to balance more steps in the strategy space.
Alternatively, matching can be done using smaller specialized LLMs trained for this task, and this would also apply to correct answer detection and path matching
8. Conclusion
In this paper, we propose using LLMs to take over the authoring of the state space for an AutoTutor-like model to minimize human time requirements and build a sample system to showcase our proposal. We find that our model performs somewhat better than a Single Prompt GPT4 model in terms of automatic metrics while maintaining full interpretability and modularity in its state space which GPT4 lacks. Our model also holds up to the desired qualities in its utterances in human evaluation, showing there is high potential in this direction. But at the same time, it still faces issues that become more apparent when going through conversations manually, indicating that there is a lot of work still to be done.
These issues can be fixed by better LLMs and more sophisticated techniques outlined in future work. Several potential steps have already been outlined in Section 4, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
Acknowledgements.
This work was funded by the Swiss National Science Foundation (Project No. 197155), and a Responsible AI grant by the Haslerstiftung. We extend our gratitude to Jakub Macina from the ETH AI Center, Nico Daheim from UKP Lab, TU Darmstadt, and Farquet Grégoire from Department of Architecture, ETH Zürich for their invaluable feedback.References
- (1)
- Anderson et al. (1985) John R. Anderson, C. Franklin Boyle, and Brian J. Reiser. 1985. Intelligent Tutoring Systems. Science 228, 4698 (1985), 456–462. https://doi.org/10.1126/science.228.4698.456 arXiv:https://www.science.org/doi/pdf/10.1126/science.228.4698.456
- Baidoo-Anu and Ansah (2023) David Baidoo-Anu and Leticia Owusu Ansah. 2023. Education in the era of generative artificial intelligence (AI): Understanding the potential benefits of ChatGPT in promoting teaching and learning. Journal of AI 7, 1 (2023), 52–62. https://dergipark.org.tr/en/pub/jai/issue/77844/1337500
- Bloom (1984) Benjamin S Bloom. 1984. The 2 sigma problem: The search for methods of group instruction as effective as one-to-one tutoring. Educational researcher 13, 6 (1984), 4–16.
- Bommasani et al. (2022) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon, Niladri Chatterji, Annie Chen, Kathleen Creel, Jared Quincy Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren Gillespie, Karan Goel, Noah Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar Khattab, Pang Wei Koh, Mark Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Ben Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, Julian Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Rob Reich, Hongyu Ren, Frieda Rong, Yusuf Roohani, Camilo Ruiz, Jack Ryan, Christopher Ré, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishnan Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2022. On the Opportunities and Risks of Foundation Models. arXiv:2108.07258 [cs.LG]
- Cai et al. (2019) Zhiqiang Cai, Xiangen Hu, and Arthur C. Graesser. 2019. Authoring Conversational Intelligent Tutoring Systems. In Adaptive Instructional Systems, Robert A. Sottilare and Jessica Schwarz (Eds.). Springer International Publishing, Cham, 593–603.
- Chen et al. (2023) Eason Chen, Ray Huang, Han-Shin Chen, Yuen-Hsien Tseng, and Liang-Yi Li. 2023. GPTutor: A ChatGPT-Powered Programming Tool for Code Explanation. In Artificial Intelligence in Education. Posters and Late Breaking Results, Workshops and Tutorials, Industry and Innovation Tracks, Practitioners, Doctoral Consortium and Blue Sky - 24th International Conference, AIED 2023, Tokyo, Japan, July 3-7, 2023, Proceedings (Communications in Computer and Information Science, Vol. 1831), Ning Wang, Genaro Rebolledo-Mendez, Vania Dimitrova, Noboru Matsuda, and Olga C. Santos (Eds.). Springer, 321–327. https://doi.org/10.1007/978-3-031-36336-8_50
- Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Alex Castro-Ros, Marie Pellat, Kevin Robinson, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Yu, Vincent Zhao, Yanping Huang, Andrew Dai, Hongkun Yu, Slav Petrov, Ed H. Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V. Le, and Jason Wei. 2022. Scaling Instruction-Finetuned Language Models. arXiv:2210.11416 [cs.LG]
- Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv preprint arXiv:2110.14168 (2021).
- Cui et al. (2023) Chenhang Cui, Yiyang Zhou, Xinyu Yang, Shirley Wu, Linjun Zhang, James Zou, and Huaxiu Yao. 2023. Holistic Analysis of Hallucination in GPT-4V(ision): Bias and Interference Challenges. arXiv:2311.03287 [cs.LG]
- Dhuliawala et al. (2023) Shehzaad Dhuliawala, Mojtaba Komeili, Jing Xu, Roberta Raileanu, Xian Li, Asli Celikyilmaz, and Jason Weston. 2023. Chain-of-Verification Reduces Hallucination in Large Language Models. arXiv:2309.11495 [cs.CL]
- Graesser et al. (2000) Arthur Graesser, Natalie Person, and Derek Harter. 2000. Teaching Tactics in AutoTutor. (01 2000).
- Graesser (2016) Arthur C. Graesser. 2016. Conversations with AutoTutor Help Students Learn. International Journal of Artificial Intelligence in Education 26, 1 (01 Mar 2016), 124–132. https://doi.org/10.1007/s40593-015-0086-4
- Graesser et al. (2001) Arthur C. Graesser, Kurt VanLehn, Carolyn P. Rose, Pamela W. Jordan, and Derek Harter. 2001. Intelligent Tutoring Systems with Conversational Dialogue. AI Magazine 22, 4 (Dec. 2001), 39. https://doi.org/10.1609/aimag.v22i4.1591
- Koedinger et al. (2004) Kenneth R. Koedinger, Vincent Aleven, Neil Heffernan, Bruce McLaren, and Matthew Hockenberry. 2004. Opening the Door to Non-programmers: Authoring Intelligent Tutor Behavior by Demonstration. In Intelligent Tutoring Systems, James C. Lester, Rosa Maria Vicari, and Fábio Paraguaçu (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 162–174.
- Kojima et al. (2023) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2023. Large Language Models are Zero-Shot Reasoners. arXiv:2205.11916 [cs.CL]
- Liu et al. (2023) Yi Liu, Gelei Deng, Zhengzi Xu, Yuekang Li, Yaowen Zheng, Ying Zhang, Lida Zhao, Tianwei Zhang, and Yang Liu. 2023. Jailbreaking ChatGPT via Prompt Engineering: An Empirical Study. arXiv:2305.13860 [cs.SE]
- Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. MathDial: A Dialogue Tutoring Dataset with Rich Pedagogical Properties Grounded in Math Reasoning Problems. arXiv:2305.14536 [cs.CL]
- Nye et al. (2014) Benjamin Nye, Arthur Graesser, and Xiangen Hu. 2014. AutoTutor and Family: A Review of 17 Years of Natural Language Tutoring. International Journal of Artificial Intelligence in Education 24 (12 2014). https://doi.org/10.1007/s40593-014-0029-5
- Nye et al. (2021) Maxwell Nye, Anders Andreassen, Guy Gur-Ari, Henryk Witold Michalewski, Jacob Austin, David Bieber, David Martin Dohan, Aitor Lewkowycz, Maarten Paul Bosma, David Luan, Charles Sutton, and Augustus Odena. 2021. Show Your Work: Scratchpads for Intermediate Computation with Language Models. https://arxiv.org/abs/2112.00114
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
- Paladines and Ramirez (2020) José Paladines and Jaime Ramirez. 2020. A Systematic Literature Review of Intelligent Tutoring Systems With Dialogue in Natural Language. IEEE Access 8 (2020), 164246–164267. https://doi.org/10.1109/ACCESS.2020.3021383
- Phung et al. (2023) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generative AI for Programming Education: Benchmarking ChatGPT, GPT-4, and Human Tutors. In Proceedings of the 2023 ACM Conference on International Computing Education Research - Volume 2 (Chicago, IL, USA) (ICER ’23). Association for Computing Machinery, New York, NY, USA, 41–42. https://doi.org/10.1145/3568812.3603476
- Rajala et al. (2023) Jaakko Rajala, Jenni Hukkanen, Maria Hartikainen, and Pia Niemelä. 2023. ”Call me Kiran” – ChatGPT as a Tutoring Chatbot in a Computer Science Course. In Proceedings of the 26th International Academic Mindtrek Conference. 83–94. https://dl.acm.org/doi/pdf/10.1145/3616961.3616974
- Reiser (2004) Brian Reiser. 2004. Scaffolding Complex Learning: The Mechanisms of Structuring and Problematizing Student Work. Journal Of The Learning Sciences 13 (01 2004), 273–304. https://doi.org/10.1207/s15327809jls1303_2
- Rich et al. (2018) Kathryn M. Rich, T. Andrew Binkowski, Carla Strickland, and Diana Franklin. 2018. Decomposition: A K-8 Computational Thinking Learning Trajectory. In Proceedings of the 2018 ACM Conference on International Computing Education Research (Espoo, Finland) (ICER ’18). Association for Computing Machinery, New York, NY, USA, 124–132. https://doi.org/10.1145/3230977.3230979
- Schmucker et al. (2023) Robin Schmucker, Meng Xia, Amos Azaria, and Tom Mitchell. 2023. Ruffle&Riley: Towards the Automated Induction of Conversational Tutoring Systems. arXiv:2310.01420 [cs.CL]
- Sleeman and Brown (1982) D. Sleeman and J.S. Brown. 1982. Intelligent Tutoring Systems. Academic Press. https://books.google.ch/books?id=pjqcAAAAMAAJ
- Song and Becker (2014) Ting Song and Kurt Becker. 2014. Expert vs. novice: Problem decomposition/recomposition in engineering design. In 2014 International Conference on interactive collaborative learning (ICL). IEEE, 181–190.
- Sonkar et al. (2023) Shashank Sonkar, Naiming Liu, Debshila Basu Mallick, and Richard G. Baraniuk. 2023. CLASS: A Design Framework for building Intelligent Tutoring Systems based on Learning Science principles. arXiv:2305.13272 [cs.CL]
- Yarkin and Coon (2022) Ferhat Yarkin and Justin P. Coon. 2022. Simple Gray Coding and LLR Calculation for MDS Modulation Systems. CoRR abs/2201.08237 (2022). arXiv:2201.08237 https://arxiv.org/abs/2201.08237
- Yu et al. (2023) Longhui Yu, Weisen Jiang, Han Shi, Jincheng Yu, Zhengying Liu, Yu Zhang, James T. Kwok, Zhenguo Li, Adrian Weller, and Weiyang Liu. 2023. MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models. arXiv:2309.12284 [cs.CL]
- Zhang et al. (2023) Muru Zhang, Ofir Press, William Merrill, Alisa Liu, and Noah A. Smith. 2023. How Language Model Hallucinations Can Snowball. arXiv:2305.13534 [cs.CL]
- Zhang et al. (2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q. Weinberger, and Yoav Artzi. 2020. BERTScore: Evaluating Text Generation with BERT. arXiv:1904.09675 [cs.CL]
| Hint | Recent history | Question | Solution Step | Rating |
Tell me this instead: What is the base number of books before Wendy adds the 20 extra books if she doubles the number of slots she has, which is 36? |
Tutor: You are correct upto where you say “Calculate the total number of slots in the shelving system by multiplying the number of rows by the number of columns. 6 rows * 6 columns = >36 slots.” What do you think should be the next step?
Wendy wants to place 20 more than double the number of books in a shelving system with 6 rows and 6 columns. How many books will she need to carry to complete her task? |
Double the number of slots to find the base number of books before adding the 20 extra books Wendy wants to place. 36 slots * 2 = <<36*2=72>>72 books. |
Grammatical:
YNN Relevant: YYY Not redundant: YYN Not revealing: YYY |
Tell me this instead: What is the profit when the total cost is subtracted from the new value of the house if the new value is $200,000 and the total cost is $130,000? |
Tutor: 200000 is correct, well done. What is the next step?
Josh decides to try flipping a house. He buys a house for $80,000 and then puts in $50,000 in repairs. This increased the value of the house by 150%. How much profit did he make? |
Find out the profit by subtracting the total cost from the new value of the house. $200,000-$130,000 = <<200000-130000=70000>>$70,000. |
Grammatical:
YYY Relevant: NNN Not redundant: NYY Not revealing: YNY |
Tell me this instead: What is the total amount of money earned in the second year from the sales of books if each was sold at the same price? |
Tutor: 65 is correct, well done. What is the next step?
Elise has been selling her Dad’s collection of 250 books for three years. Each book sells at 20$, and she sold twice as many books in the first year as she has sold in the current year. There are currently 50 unsold books, and her sales number this year is 45. What’s the total amount of money she earned in the second year? |
Compute the total amount of money earned in the second year by multiplying the number of books sold by the price per book: <<65 * 20 = 1300>>1300 dollars. |
Grammatical:
YYY Relevant: YYY Not redundant: NNN Not revealing: YYN |
Tell me this instead: What is the total cost per person for engaging in minigolf, purchasing arcade tokens, and enjoying two go-kart rides? |
Tutor:What should be the first step?
Morgan’s dad said that she had $90 budgeted for her birthday party. She wants to make sure she and her friends all get to play one round of mini-golf, have $5 in arcade tokens, and get to ride the go-karts twice. A round of mini-golf is $5. The Go-karts cost $10 a ride. How many friends can she invite? |
Calculate the total cost per person for mini-golf, arcade tokens, and two go-kart rides. 5 + 5 + (10 x 2) = <<5 + 5 + (10 * 2)=30>>30 dollars per person. |
Grammatical:
YYY Relevant: YYY Not redundant: YYY Not revealing: NYN |
| Hint | Recent history | Question | Solution Step | Rating |
That is not quite right. Here is a Hint: To calculate the total monthly expense for Madeline’s dog, you need to sum the individual costs for food, treats, and medicine. Consider each expense as a separate amount and add them together to find the total. Remember, the food expense is a two-digit number ending in zero, and both the treat and medicine expenses have the same amount, which is also a two-digit number ending in zero |
Tutor: Tell me this instead: What is the total monthly expense for Madeline’s dog when combining the costs of food, treats, and medicine?
Every month, Madeline has to buy food, treats, and medicine for her dog. Food costs $25 per week. Treats cost $20 per month. Medicine costs $100 per month. How much money does Madeline spend on her dog per year if there are 4 weeks in a month? |
Add the monthly costs of food, treats, and medicine to find the total monthly expenses for Madeline’s dog. $100 (food per month)+ $20 (treats per month) +$100 (medicine per month) = 100+20+100=220$220 per month |
Grammatical:
YNY Specific: YNN Helpful: YNY Not revealing: YYY |
That is not quite right. Here is a Hint: To determine the number of games Bobby has, first identify Brian’s games and multiply that amount by three. Once you have that number, decrease it by five and you will find out Bobby’s total games. |
Tutor: Tell me this instead: What is the number of games Bobby has if you take away 5 from three times the number of Brian’s games?
Brian’s friend Bobby has 5 fewer than 3 times as many video games as Brian does. If Brian has 20 video games but lost 5 right before the comparison was made, how many does Bobby have? |
Subtract 5 from the tripled number of Brian’s games to find out how many games Bobby has.
Grammatical:
YYY Specific: YYY Helpful: YYY Not revealing: NNY |
An example solution and utterance space from the MultiArith dataset obtained using GPT4